OpenStudy (anonymous):

The gradient to the curve y=2x^2+x-1 at point A is the perpendicular to the straight line 5y=-x. find coordinate of A.

OpenStudy (anonymous):

gradient in this case means slope, right ? find the slope of the straight line first, \[5y = -x\]\[y = -\frac{1}{5}x\]\[m_{1} = -1/5\]because the slope at point A is perpendicular to the slope of the straight line,\[m_{1}m_{2} = -1\]\[m_{2} = 5\]then we can find point A by using the first derivative of the curve\[y = 2x^2 + x - 1\]\[y' = 4x + 1\]\[5 = 4x + 1\]\[x = 1\]\[y = 2(1)^2 + 1 - 1\]\[y = 2\]point A is (1,2)

OpenStudy (anonymous):

How to know when to use derivative??

OpenStudy (anonymous):

y′=4x+1 5=4x+1<< i dont understand this part?? where do u get the five. x=1

OpenStudy (anonymous):

to find the slope of a curve at a certain point you need to use the first derivative of the curve

OpenStudy (anonymous):

\[y' = 4x + 1\]\[m_{2} = 4x + 1\]

OpenStudy (anonymous):

TQ. Very much!! From that, how to find the equation of the tangent to the curve at point A?

OpenStudy (anonymous):

recall the equation of a line\[y - y_{1} = m(x - x_{1})\]plug (1,2) for x1 and y1, and 5 for m

OpenStudy (anonymous):

Is it y=5x-3 ? TQ very much.